Kyle inherits $40,000 and invests part of the money in bonds with an interest rate of 11 percent. The remainder of the money is in savings at a 9 percent rate. What amount has he invested at each rate if he receives $4,080 in interest for one year?

Answer by mananth(16946)

Bonds 11.00% per annum ----x
Savings 6.00% per annum ----y
The sum of the investments is $40,000.00
The sum of individual interests = $4,080.00
x + y 40000 ------------------------1
11.00% x + 6.00% y = $4,080.00
Multiply by 100
11 x + 6 y = $408,000.00 --------2

Multiply (1) by -11
we get

-11 x -11 y = -440000.00

Add this to (2)

0 x -5 y = -$32,000.00

divide by -5

y = $6,400.00 investment at 6.00%
Balance $33,600.00 investment at 11.00%
CHECK
$33,600.00 @ 11.00% $3,696.00
$6,400.00 @ 6.00% $384.00
Total --------------- $4,080.00